A rocket beigns from rest and accelerates upward at +0.5m/s^2.
a) At what time will the rocket reach a height of 1000 m?
b) What will be the rocket's velocity at that point?

Answer by Alan3354(69443)

A rocket beigns from rest and accelerates upward at +0.5m/s^2.
a) At what time will the rocket reach a height of 1000 m?
h(t) = at^2/2
at^2/2 = 1000
0.5*t^2 = 1000
t^2 = 2000
t = sqrt(2000) seconds
===============
b) What will be the rocket's velocity at that point?
v = a*t = 0.5*sqrt(2000) m/sec